US national health expenditure was $7,892 per person in 2008 and $10,364 in 2016. The Consumer Price Index had a value of 210.228 in 2008 and a value of 241.432 in 2016. In 2016 dollars, how much was spending in 2008?
Answer #1
Year Expenditure CPI
2008 $ 7,892 210.228
2016 $ 10,364 241.432

Spending in the year 2008 has to be converted into 2016 dollar.

\large Amount = \$ 7,892 \times \frac{CPI_{2016}}{CPI_{2008}}

\large \implies Amount = \$ 7,892 \times \frac{241.432}{210.228}

\large \implies Amount = \$ 7,892 \times 1.148429

\large \implies Amount \approx \$ \ 9,063.40

Spending in 2008 in the 2016 dollar = $ 9,063

Please contact if having any query will be obliged to you for your generous support. Your help mean a lot to me, please help. Thank you.